COHERENT RAMAN SPECTRA OF THE ν_{1} MODE OF ^{10}BF_{3} AND ^{11}BF_{3}

\underline{ROBYNNE KIRKPATRICK}, JOSEPH W. NIBLER, Department of Chemistry, Oregon State University, 153 Gilbert Hall, Corvallis, OR 97331-4003; TONY MASIELLO, Pacific Northwest National Laboratory, P. O. Box 999, Mail Stop K8-88, Richland, WA 99532; ALFONS WEBER, National Institute of Standards and Technology, Gaithersburg, MD 20899.

High resolution (0.001 cm^{−1}) coherent anti-Stokes Raman spectroscopy (CARS) was used to directly examine the ν_{1} symmetric stretching mode of the D_{3h} molecules ^{10}BF_{3} and ^{11}BF_{3}. Calculated spectra were constructed using rovibrational parameters deduced from published infrared hot band, combination band, and difference band studies, and the similarity to the experimental CARS spectra confirms the validity of the constants. No significant perturbations by Fermi resonance or Coriolis interactions with nearby states are observed, in marked contrast to the case of sulfur trioxide, a similar D_{3h} molecule recently studied. Since the boron nucleus is at the center of mass, the ^{10}BF_{3} and ^{11}BF_{3} ν_{1} Q-branches are the same according to the harmonic oscillator model. Interestingly, the observed frequency of ^{11}BF_{3} is 0.198 cm^{−1} higher than that of ^{10}BF_{3}. This result is reproduced almost exactly (0.200 cm^{−1}) using ab initio calculations (B3LYP/cc-pVTZ) that included evaluation of cubic and quartic force constants and x_{ij} anharmonicity constants. Ab initio methods also predict to within 1% the ΔB and ΔC changes in the rotational constants in going from the ground state to the v = 1 vibrational level.